OpenStudy (anonymous):

The function g(x) is obtained by translating f(x) = 2^x to the right 6 units and down 3 units. Write an equation for g(x).

OpenStudy (cwrw238):

f(x) ---- > f(x - 6) = move to right 6 places f(x) ------> f(x) - 3 = move 3 units down

OpenStudy (anonymous):

wait so is that how you write the equation

OpenStudy (cwrw238):

g(x) = 2^(x - 6) - 3

OpenStudy (anonymous):

oook thanks

OpenStudy (cwrw238):

x-6 replaces x -3 does the 3 moves down